The following transactions pertain to 2018, the first-year operations of Adams Company. All inventory was started and completed during 2018. Assume that all transactions are cash transactions.

  1. Acquired $4,300 cash by issuing common stock.

  2. Paid $710 for materials used to produce inventory.

  3. Paid $1,890 to production workers.

  4. Paid $520 rental fee for production equipment.

  5. Paid $100 to administrative employees.

  6. Paid $114 rental fee for administrative office equipment.

  7. Produced 300 units of inventory of which 210 units were sold at a price of $13 each.

Complete this question by entering your answer in the tabs below Income Statement Balance Sheet Prepare an income statement in accordance with GAAP. (Do not round your intermediate calculations.) ADAMS COMPANY Income Statement for 2018 < Income Statement Balance Sheet >

Complete this question by entering your answer in the tabs below. Income Statement Balance Sheet Prepare a balance sheet in accordance with GAAP. (Do not round your intermediate calculations.) ADAMS COMPANY Balance Sheet as of 12131/2018 Assets Total assets Equity Total equity K Income Statement Balance Sheet>
Answer #1
INCOME STATEMENT
Sales (210*$13) 2730
Cost of goods sold (see workings below) 2184
Gross profit 546
Operating expenses:
Salary to administrative employees 100
Rental fee for administrative office equipment 114
Net income 332
WORKINGS:
Cost of goods sold:
Cost of production = 710+1890+520 = 3120
Cost of ending inventory = 3120*90/300 = 936
COGS 2184
BALANCE SHEET
ASSETS:
Cash (4300-710-1890-520-100-114+2730) 3696
Inventory 936
Total assets 4632
EQUITY:
Common stock 4300
Retained earnings 332
Total liabilities and equity 4632